OpenStudy (anonymous):

y = x2 - 5 write in vertex form

OpenStudy (anonymous):

is x2 really x^2?

OpenStudy (anonymous):

x^2 sorry about that

OpenStudy (anonymous):

it's ok ...i think it's already in vertex form actually

OpenStudy (anonymous):

it cant be the answers are a.y+(x+5)^2-5 b.y=(x-0)^2-5 c.y=(x-5)^2+0 d.y=(x+5)^2-0

OpenStudy (anonymous):

b is the answer it's the same equation

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can you help me with this one y = x^2 + 2x + 1 - 1 vertex form

OpenStudy (anonymous):

a.y=(x-1)^2-1 b.y=(x+2)^2-1 c.y=(x+1)^2-1 D.y=(x-2)^2-1

OpenStudy (anonymous):

I think you have to complete the square for this one. Do you mean to have 1-1 at the end of the equation?